OpenStudy (katherinesmith):

What is the equation of a parabola with a vertex at (0, 0) and a focus at (0, 6)

OpenStudy (anonymous):

do you know what it looks like approximately?

OpenStudy (katherinesmith):

not at all.

OpenStudy (anonymous):

well that was wrong lets try again

OpenStudy (anonymous):



OpenStudy (katherinesmith):

i know what the graph looks like. i meant i didn't know what the equation looks like. haha

OpenStudy (anonymous):

vertex is at the origin \((0,0)\) and focus is at \((0,6)\)which is 6 units up we need to know what it looks like so we know that the \(x\) term is squared, i.e. it the equation will look like \[4py=x^2\]

OpenStudy (anonymous):

and in your case \(p=6\) so you get \[24y=x^2\]

OpenStudy (katherinesmith):

now i have to break down the x^2 to be x =

OpenStudy (anonymous):

well no, there is no breaking down \(x^2\) whatever that means you cannot write this as \(x=\text{ something }\)

OpenStudy (katherinesmith):

are you sure? okay.

OpenStudy (anonymous):

i am sure the \(x\) is squared, it is \(x^2\) not \(x\)

OpenStudy (anonymous):

lets check it http://www.wolframalpha.com/input/?i=parabola+24y%3Dx^2